You are the finance director of the Australian listed company, Fire Ltd that has a functional currency in AUD. Fire Ltd that has borrowings from a US bank. The company's financial year ends on 30 June 2020. Fire Ltd entered the following transaction during the year.

On 1 January 2020, Fire Ltd borrowed US $6,000,000 from an investment bank in the United States for a 12-month period. The borrowing has a fixed rate of interest at 10% p.a. payable at 6- month intervals. On 1 April 2020, Fire Ltd entered a 9- month forward contract to buy US $ 5,000,000 in order to hedge against the foreign exchange risk on the US $ loan principal. The following exchange rates apply.

Date Spot rate Forward rate for 31/12/2020
1 Jan 2020 A$1 = US$ 0.89 A$1 = US$ 0.84
1 Apr 2020 A$1 = US $ 0.86 A$1 = US$ 0.82
30 June 2020 A$1 = US $ 0.85 A$1 = US $ 0.79
31 Dec 2020 A$1 = US $ 0.80

A$1 = US $ 0.80

Assume 0% discount rate for fair value calculations.

Prepare the entries of Fire Ltd to account for its foreign currency transactions.
Answer #1

1.) On 1st January 2020 Fire Ltd, (the company) borrows US $ 6,000,000 for a period of 12 months at the exchange rate of US$/AS$ 0.89

Therefore AS$ equivalent borrowing = US$ 6,000,000/.89 = 6,741,573.03

Entry : Dr Bank A/c 6,741,573.03

Cr Loan A/c 6,741,573.03

2.) On 30th June 2020 the company is liable to pay interest towards loan after 6 months

Interest liability = US $ 6,000,000*10%*6/12 = US$ 300,000.

This will be converted at the spot rate as 30th June 2020 ie. US$/AS$ 0.85

AS$ equivalent interest = US$ 300,000/.85 = AS$ 352,941.18

Entry : Dr Interest on loan A/c 352,941.18

Cr Bank A/c 352,941.18

3.) On 31st December 2020 we have two transactions

i) Interest payment

Interest liability = US $ 6,000,000*10%*6/12 = US$ 300,000.

This will be converted at the spot rate as on 31st December 2020 ie.US$/AS$ 0.80

AS$ equivalent interest = US$ 300,000/.80 = AS$ 375,000

Entry : Dr Interest on loan A/c 375,000

Cr Bank A/c 375,000

ii) Principal repayment

The company has taken a forward contract for repayment of principle on 1st April 2020 at US$/AS$ .082 for US$ 5,000,000.Whereas the loan payable is US$ 600,000

Therefore principal settlement will happen in two parts

a) US$ 500,000 will be settled at the rate of US$/AS$ 0.82

AS$ equivalent amount = US$ 5,000,000/0.82 = 6,097,560.97

b) US$ 100,000 will be settled at spot rate of US$/AS$ 0.80

AS$ equivalent amount = US$ 1,000,000/0.80 = 1,250,000

Therefore total payment towards loan in AS$ = 6,097,560.97+1,250,000 =7,347,560.97

Increase in payment towards loan due to exchange rate fluctuation = 7,347,560.97-6,741,573.03 = 6,05,987.94 (which is a loss to the company)

Entry : Dr Loan A/c 6,741,573.03

Dr Exchange fluctuation loss A/c 6,05,987.94

Cr Bank A/c 7,347,560.97

Note : There will be no entry for purchase of forward contract unless there is any upfront premium required to be paid